A right-angle triangle has a height 8cm more than twice the length of the base. If the area of the triangle is 96cm, find the demensions of the triangle. Answer by stanbon(75887) 
 A right-angle triangle has a height 8cm more than twice the length of the base. If the area of the triangle is 96cm, find the demensions of the triangle.
-------------
Area = (1/2)(base)(height)
96 cm^2 = (1/2)(x)(2x+8)
96 = x^2+4x
x^2+4x-96=0
(x+12)(x-8)=0
Positive solution: x=8 cm (length of the base)
2x+8 = 24 cm (the height of the triangle)
==========
Cheers,
Stan H.